# RAP: 3D Rasterization Augmented End-to-End Planning

### v1

@@ 1 Introduction
We equip RAP with multiple planners, showing that its benefits hold across diverse model designs. Extensive experiments show that RAP achieves strong closed-loop robustness and long-tail generalization, ranking 1st place on the following 4 major E2E planning benchmarks: NAVSIM v1/v2 (Dauner et al., 2024; Cao et al., 2025), Waymo Open Dataset (WOD) Vision-based E2E Driving (Ettinger et al., 2021), and Bench2Drive (Jia et al., 2024). We find that efficient rasterization with feature-space alignment provides the scalability and robustness needed for end-to-end planning, without requiring photorealistic rendering.
@@ 2 Related Work
All primitives are rasterized into an RGB canvas $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ using depth-aware compositing, where each fragment stores depth $d$ with a fading weight $\alpha=\max(0,\,1-d/d_{\max})$ blended through a single buffer to resolve occlusions; primitives crossing the view boundary are clipped with the Sutherland–Hodgman (Sutherland & Hodgman, 1974) operator to ensure stable visibility. Focusing on geometric and semantic fidelity while discarding photorealistic details, our rasterizer delivers spatial cues essential for planning at a fraction of the cost, enabling large-scale counterfactual generation. As shown in Figure 4, features extracted by a frozen DINOv3 encoder and visualized via PCA remain qualitatively consistent between rasterized and real images, indicating that abstraction preserves perceptual cues crucial for downstream learning. Figure 3 shows a qualitative comparison between real and rasterized views across a 5 s horizon.
@@ 3.3 Raster-to-Real Alignment
Global alignment complements spatial-level supervision by keeping the overall feature distributions consistent. For instance, rasterized views may contain pure black backgrounds absent in real data; aligning globally mitigates such biases and improves generalization. Besides, since rasterized samples vastly outnumber real–raster pairs, unsupervised domain adaptation (Ganin & Lempitsky, 2015) provides an effective way to exploit the full synthetic corpus by enforcing global consistency even without paired supervision.
For each sample, we compute a global representation $g\in\mathbb{R}^{d^{\prime}}$ by average pooling its feature map $F\in\mathbb{R}^{N\times d^{\prime}}$. We then train a domain classifier $D$ to predict whether $g$ comes from real or rasterized data. Following Ganin & Lempitsky (2015), a gradient reversal layer is inserted before $D$, such that the encoder is optimized to maximize domain confusion while $D$ is optimized to minimize classification error with $y\in\{0,1\}$ the domain label:
@@ 4.1 Leaderboard Evaluation
Table 3: Top-6 entries on the public leaderboard for the WOD Vision-based E2E Driving Challenge (up to September 2025).
@@ 5 Conclusion
We presented Rasterization Augmented Planning (RAP), a scalable framework for end-to-end driving that leverages lightweight 3D rasterization and feature-space alignment as an alternative to costly photorealistic rendering. By enabling recovery-oriented perturbations and cross-agent synthesis, RAP scales training to large-scale counterfactual scenes while preserving semantic and geometric fidelity. Extensive experiments across four benchmarks demonstrate that RAP consistently improves closed-loop robustness and long-tail generalization, establishing a practical and effective recipe for scaling end-to-end autonomous driving. Limitations & Future Work: a key limitation of our approach is that it remains within the imitation learning paradigm, which inherits issues such as causal confusion. In future work, we aim to extend 3D rasterization into a full simulator to support closed-loop reinforcement learning, enabling richer interaction and policy improvement beyond offline demonstrations.
@@ Appendix A Appendix
A.1 LLM Usage
@@
Optimizer AdamW (Loshchilov & Hutter, 2017)
@@
A.2 More Training Details
@@
For adversarial global alignment, we adopt a gradient reversal layer following Ganin & Lempitsky (2015). During the forward pass, the GRL acts as an identity mapping, while in the backward pass it multiplies the gradients by $\lambda$, forcing the encoder to learn domain-invariant features. We schedule $\lambda$ with a smooth annealing function
@@
WOD Vision-based E2E Driving Finetuning.
For finetuning on the Waymo Open Dataset (WOD) Vision-based E2E Driving benchmark, we use a learning rate of $1\times 10^{-5}$ and unfreeze the visual encoder. Finetuning is performed in two stages: first on the official training split, and then on the validation split with a batch size of 16. For the final test submission, we apply non-maximum suppression (NMS) ensembling over two checkpoints to improve robustness.
Bench2Drive Mixed-Training and Dataset Alignment.
A.3 Visualizations for ablation study on 3D Rasterization

### v2

@@ 1 Introduction
We equip RAP with multiple planners, showing that its benefits hold across diverse model designs. Extensive experiments show that RAP achieves strong closed-loop robustness and long-tail generalization, ranking 1st place on the following 4 major E2E planning benchmarks: NAVSIM v1/v2 (Dauner et al., 2024; Cao et al., 2025), Waymo Open Dataset Vision-based E2E Driving (WOD-E2E) (Xu et al., 2025), and Bench2Drive (Jia et al., 2024). We find that efficient rasterization with feature-space alignment provides the scalability and robustness needed for end-to-end planning, without requiring photorealistic rendering.
@@ 2 Related Work
All primitives are rasterized into an RGB canvas $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ using depth-aware compositing, where each fragment stores depth $d$ with a fading weight $\alpha=\max(0,\,1-d/d_{\max})$ blended through a single buffer to resolve occlusions; primitives crossing the view boundary are clipped with the Sutherland–Hodgman (Sutherland and Hodgman, 1974) operator to ensure stable visibility. Focusing on geometric and semantic fidelity while discarding photorealistic details, our rasterizer delivers spatial cues essential for planning at a fraction of the cost, enabling large-scale counterfactual generation. As shown in Figure 4, features extracted by a frozen DINOv3 encoder and visualized via PCA remain qualitatively consistent between rasterized and real images, indicating that abstraction preserves perceptual cues crucial for downstream learning. Figure 3 shows a qualitative comparison between real and rasterized views across a 5 s horizon.
@@ 3.3 Raster-to-Real Alignment
Global alignment complements spatial-level supervision by keeping the overall feature distributions consistent. For instance, rasterized views may contain pure black backgrounds absent in real data; aligning globally mitigates such biases and improves generalization. Besides, since rasterized samples vastly outnumber real–raster pairs, unsupervised domain adaptation (Ganin and Lempitsky, 2015) provides an effective way to exploit the full synthetic corpus by enforcing global consistency even without paired supervision.
For each sample, we compute a global representation $g\in\mathbb{R}^{d^{\prime}}$ by average pooling its feature map $F\in\mathbb{R}^{N\times d^{\prime}}$. We then train a domain classifier $D$ to predict whether $g$ comes from real or rasterized data. Following Ganin and Lempitsky (2015), a gradient reversal layer is inserted before $D$, such that the encoder is optimized to maximize domain confusion while $D$ is optimized to minimize classification error with $y\in\{0,1\}$ the domain label:
@@ 4.1 Leaderboard Evaluation
Table 3: Top-6 entries on the public leaderboard for the WOD-E2E Driving Challenge (up to September 2025).
@@ 5 Conclusion
We presented RAP, a scalable framework for end-to-end driving that leverages lightweight 3D rasterization and feature-space alignment as an alternative to costly photorealistic rendering. By enabling recovery-oriented perturbations and cross-agent synthesis, RAP scales training to large-scale counterfactual scenes while preserving semantic and geometric fidelity. Extensive experiments across four benchmarks demonstrate that RAP consistently improves closed-loop robustness and long-tail generalization, establishing a practical and effective recipe for scaling end-to-end autonomous driving. Limitations & Future Work: a key limitation of our approach is that it remains within the imitation learning paradigm, which inherits issues such as causal confusion. In future work, we aim to extend 3D rasterization into a full simulator to support closed-loop reinforcement learning, enabling richer interaction and policy improvement beyond offline demonstrations.
@@ Acknowledgments
Ethics Statement
This work adheres to the ICLR Code of Ethics. Our research focuses on developing scalable training methods for end-to-end autonomous driving planners using synthetic rasterized data. No human subjects were involved, and all datasets used (nuPlan, Waymo, NAVSIM, and Bench2Drive) are publicly released under their respective licenses. We acknowledge that autonomous driving research raises safety and fairness concerns. By emphasizing reproducibility, transparent methodology, and evaluation on diverse benchmarks, we aim to mitigate risks of overclaiming model capabilities. The proposed method does not introduce new privacy, security, or discrimination risks beyond those already present in the underlying datasets.
Reproducibility Statement
We have made efforts to ensure that our work is reproducible. The paper provides detailed descriptions of the model architecture (Section 3), training setup (Section 4), and ablation studies (Section 4.2). Additional hyperparameters and implementation details are provided in the Appendix. All datasets used are publicly available (nuPlan, WOD, Bench2Drive), and we describe the curation and filtering process in Section 4. Source code and configuration files will be submitted and released with the camera-ready version to facilitate full reproducibility. These instructions apply to everyone, regardless of the formatter being used.
@@ Appendix A Appendix
A.1 Discussions
Does Simplified 3D Rasterization Risk Overlook Critical Real-World Visual Cues?
Our 3D rasterization pipeline simplifies appearance by rendering only the annotated geometric elements of the scene, such as lanes, vehicles, and traffic signals. This choice reflects a deliberate trade-off. Rasterization gives us the speed and controllability needed to generate large-scale, diverse training data. High-fidelity rendering methods cannot reach this scale. At the same time, rasterization does not aim to replace real images. It complements them.
Despite this simplification, we find that the model retains strong perception and planning ability when applied to real-world images. Qualitative results on nuPlan (Figure 7) show that RAP correctly reacts to visual elements that are not present in the rasterization ontology, such as unannotated “Keep Left” signs or LED displays on a truck. Additional long-tail cases from WOD-E2E (Figure 8) further confirm that the model can handle subtle and unexpected signals at inference time.
These results indicate that simplified 3D rasterization does not hinder the model’s ability to reason about fine-grained or OOD cues. Instead, it provides a strong geometric foundation that supports robust, scalable learning. Real images, seen both during training and always at inference, supply the remaining richness that rasterization cannot encode. The strong performance of RAP on the corner-case-focused WOD-E2E benchmark (Table 3) reinforces this conclusion.
Does Real-to-Sim Feature Alignment Cause Information Loss?
A natural concern is that aligning real-world features to simplified rasterized features may suppress unannotated or subtle visual cues. Our results show that this does not occur, largely due to the multi-task learning design of our planner. The Real-to-Raster alignment provides a stable geometric prior, while the planning and perception objectives applied on real features compel the model to retain all task-relevant information necessary for accurate decision-making in complex traffic scenes.
Table 7: Comparison of alignment directions under 50% real data. Lower ADE is better.
[…]
Raster-to-Real 1.12
Symmetric Alignment 1.14
Real-to-Raster (ours) 1.02
To further understand this effect, we compare three alignment strategies: Raster-to-Real, symmetric alignment, and our Real-to-Raster variant. Following the same ablation setup in Section 4.2, we fix the real-data ratio to 50%. The results in Table 7 show that Real-to-Raster yields the lowest ADE, outperforming both alternatives. This supports our hypothesis that raster features serve as a clean structural scaffold, while the planning loss ensures that semantic richness from the real image is preserved. These findings align with recent representation alignment literature (Yu et al., 2024), where aligning to a well-structured feature space improves semantic abstraction without discarding perceptual detail. Future work may explore adaptive or task-aware alignment schemes to further reduce potential information loss.
Figure 7: RAP-DINO retains fidelity to fine-grained real-world cues. This figure ablates our single, fully-trained model by comparing its predictions (red trajectory) against the ground truth (green trajectory) when conditioned on two different inputs: the raw real-world image (Top) vs. the simplified rasterized image (Bottom). (Left) Scenario A: An unannotated “Keep Left” sign, absent from the raster data, is correctly perceived from the raw image (Top), leading to a correct trajectory. The raster-conditioned model (Bottom) fails. (Right) Scenario B: A challenging OOD dynamic LED arrow, also absent from the raster, is successfully identified from the raw image (Top), resulting in a safe lane change.
Figure 8: Qualitative Results on the WOD Vision-Based E2E Driving Dataset. The red trajectory shows the predictions of RAP-DINO, and the green trajectory denotes the ground-truth future path. Across the four examples, RAP-DINO successfully identifies a stop sign, traffic cones, and oncoming vehicles at night, leading to accurate and safe trajectory plans.
WOD Vision-based E2E Driving Finetuning.
For finetuning on the WOD-E2E Driving benchmark, we use a learning rate of $1\times 10^{-5}$ and unfreeze the visual encoder. Finetuning is performed in two stages: first on the official training split, and then on the validation split with a batch size of 16. For the final test submission, we apply non-maximum suppression (NMS) ensembling over two checkpoints to improve robustness.
Bench2Drive Mixed-Training and Dataset Alignment.
A.2 Visualizations for ablation study on 3D Rasterization
A.3 LLM Usage
@@
Optimizer AdamW (Loshchilov and Hutter, 2017)
@@
A.4 More Training Details
@@
For adversarial global alignment, we adopt a gradient reversal layer following Ganin and Lempitsky (2015). During the forward pass, the GRL acts as an identity mapping, while in the backward pass it multiplies the gradients by $\lambda$, forcing the encoder to learn domain-invariant features. We schedule $\lambda$ with a smooth annealing function